# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 5 Enhancing Recurrent Neural Networks
This section gathers recent approaches that are not directly related to the previous macro-categories of Transformers architectures and State-Space Models, but still focus on improving recurrent models. It turns out that several of the architectural trends described in the previous sections (i.e., element-wise linear recurrence, novel gating mechanisms, etc.), and other new ones, are also significantly explored in the scientific literature that aims at facing two of the main drawbacks of RNN: slow sequential training and limited capability in modeling long-term dependencies. In Figure 16 we report an overview of the main topics/approaches covered by this Section.
Figure 16: Overview of the organization of Section 5.
Simplifying RNNs to Gain Speed.   RNNs are based on sequential processing of the input data, which does not directly allow to build efficient implementations that process the input tokens in parallel or that update the components of the hidden state in parallel. It turns out that this limit is mostly due to ($i$) the non-linearity applied to recurrent layers and to ($ii$) the fact that updates in the hidden state are performed by full matrix multiplication, due to a dependency on all components of the hidden states from the previous time step. In detail, the standard update scheme of RNNs (Eq. 2) assumes that all the neurons in one layer contribute to the state computation of every other neuron (i.e., through the $Ax_{t-1}$ term). Each element of the state vector $x_{t}$ depends on all the entries of $x_{t-1}$. Early works, such as Independently RNN (IndRNN) [157], propose layers composed by “independent” neurons, achieved by modifying Eq. (2) as,
@@
where Eqs. (57)-(58) represent the proposed “lightweight” recurrence, where the reset gate $r_{t}$ adaptively combines the input and the cell state $c_{t}$, with learnable vectors $a_{f},a_{r},b_{r},b_{f}$ and learnable matrices $B_{f},B_{r}$. The skip connection in Eq. (60) favours gradient propagation. Independence between distinct hidden states enables efficient element-wise product instead of full matrix multiplication (i.e., in classical forget gates dense matrices products inject a dependencies on all neurons previous states), as the authors show when the (nonlinear) recurrence is fused within a single CUDA kernel. This allows to reduce the complexity to $\mathcal{O}(Lb{d_{\text{in}}})$, where $b$ denotes here the batch dimension, while a standard LSTM takes $\mathcal{O}(Lb{d_{\text{in}}}^{2})$. The seminal work by Martin & Cundy [185] highlights that linear recurrences in the form of $x_{t}=\Lambda_{t}x_{t-1}+u_{t}$ are specific instances of the scan operation, a computation involving the repeated application of a binary operator over an array of data. This allows for a highly parallelizable unrolling of the recurrence using parallel scans [23], resulting in substantial improvements in training speeds. When $\Lambda_{t}$ is diagonal, the cost of a ELR with parallel scan and $p$ processors is $\mathcal{O}\left(6{d_{\text{s}}}(L/p+\text{log}p)\right)$, while the cost of serial scan is $\mathcal{O}(2{d_{\text{s}}}L)$. This reduction becomes important when the sequence length $L$ is large since, given sufficient processors $p$, the parallel time scales logarithmically with the sequence length. Please refer to Appendix H of [271] for a detailed overview of the parallel scan operation. Moreover, several existing models (QRNN[25], SRU [153]) fall under this class of approaches, and the authors of [185] provide an efficient CUDA kernel that speed up their training. This work laid the foundations for the efficient parallel implementation of SSMs such as S5 [271] and others [213] (Section 4). Additionally, while typical forget gate values depend on both the previous hidden state and the current input, the authors suggest that forget gate values should depend solely on the current inputs to enable parallel training. Lately, Lim et al. [165] built on top of [185] and showed that it is also possible to parallelize the evaluation and training of non-linear sequential models like classic RNN and NeuralODE [38], by introducing a general framework to solve non-linear differential equations, which are restated as fixed-point iteration problems with quadratic convergence, equivalent to Newton’s method. Each fixed-point iteration involves parallelizable operations and an inverse linear operator that can be evaluated in parallel even for the aforementioned sequential models, resulting in improvements up to $3$ orders of magnitude in terms of speed in the case of long sequences.
@@
Enhancing Gating Mechanisms.   Learning long term dependencies requires the ability to modulate the effect of the incoming information. Several recent studies have incorporated gating mechanisms inspired by LSTMs [105] (or related intuitions) into SSMs, which are characterized by gates acting on linear recurrence layers [89, 85, 271], yielding impressive performance gains. Gu et al. [89] investigated the saturation regime of gates, remarking the fact that capturing long-term dependencies in gated RNNs requires forget gate values close to one. Unfortunately, learning with gates in their saturated regimes (i.e., values close to 00 or $1$) is difficult, since they suffer from vanishing gradients. Moreover, if all forget gate values are close to one, the model ability to forget irrelevant information is hindered. To overcome such issues, the authors of [89] propose to tweak the forget gate $f_{t}$ with an independent refine gate $r_{t}$ that is exploited to produce an input-dependent bounded additive term $\phi(f_{t},u_{t})$ that modulates $f_{t}$, allowing much higher/lower activations,
@@
where $W_{r}^{*}$ denote projection matrices and $\hat{c}_{t}$ is the cell input activation vector from LSTMs. The form of the additive update $\phi(f_{t},u)$ emerges from the consideration that gating mechanisms should be bounded in $[0,1]$, symmetric around 00 and differentiable. Additionally, the authors propose to initialize the forget gate $f_{t}$ with values sampled from the uniform distribution $\mathcal{U}(0,1)$, instead of constant values, even allowing negative biases. This choice foster the gate ability to grasp different timescales as noticed in the chrono initialization approach by Tallec & Olivier [280]. In previous Sections, we showed how several linear RNNs with static decay rates perform eigendecompositions on the recurrent weight matrix to achieve element-wise linear recurrence [111, 213]. Notice that, if only real-valued eigenvalues are allowed, this choice restricts the range of the recurrent weight matrix to be symmetric, limiting the expressiveness of the model. To overcome this limitation, linear RNNs often employ complex-valued eigenvalues [90, 213]. Following such intuitions, Hierarchically Gated Recurrent Units (HGRU) [235] exploits linear recurrence in the complex domain, and addresses the saturating issue pointed out by Gu et al. [89] by adding an additive learnable value $\Gamma$ to the original forget gate, with the purpose of pushing gate activations away from the saturated regimes. The $\Gamma$ variable, which acts as a lower bound on forget gate values, is forced to increase monotonically with the model depth, inspired from Ordered Neuron LSTM [265]: small value in lower layers, in order to ease the forgetting of past-information (short-term dependencies); forget value close to one in top-most layers, facilitating the modeling of long-term dependencies. In details, HGRU leverages a gated linear recurrent as follows,
@@
where $\Gamma\in\mathbb{R}^{H\times{d_{\text{s}}}}$ independently parametrizes the lower bounds for all hidden states, where cumsum and softmax operate over the first dimension of their tensor input. Basically, the composition of softmax and cumsum form an activation function which yields a monotonically increasing vector in $[0,1]$. The squared bracket-based notation is defined as $[\text{cumsum}(x)]_{l}=(\sum_{i=1}^{l}x_{i})-x_{1}$. Notice that $\text{cumsum}_{0}$ is applied to the layer dimension across different layers, to enable upper layers in modeling long-range dependencies. Then, the model exploits an output projection with a learned gate, similarly to what happens in SSMs. Another perspective on the role of gating mechanisms can be appreciated by inspecting the already described connection between the update mechanisms in linear RNNs and linear Transformers, of the previous Sections. Recently, Zucchet et al. [319] proposed a unifying view on such architectures, driven by the role of gating functions. In particular, under a specific parametrization (that leverages GLU [56] and requires a number of parameters squared with respect to attention parameters), they showed that RNNs equipped with linear recurrent layers interconnected by feed-forward paths with multiplicative gating can, through learning, encode attention-based algorithms disguised in their weights.
@@
where $q_{t},k_{t},v_{t}\in\mathbb{R}^{d_{\text{k}}}$ are linear projections of the input (such as in self-attention), while $w_{i},w_{f}\in\mathbb{R}^{{d_{\text{in}}}}$ are learnable weight vectors. In the cell-state update rule, the forget gate acts as a decay rate, the input gate as a learning rate, while the output gate scales the vector which is retrieved by the outer product. The normalizer state $n_{t}$, which keeps a record of the “strength” of the gates, is the weighted sum of the key vectors, where, in turn, each key is weighted by the input gate and the current forget gate. In mLSTM, considering multiple cells is equivalent to multiple heads, since in this case there is no memory mixing, due to the presence of matrix states. Interestingly, the absence of memory mixing (no hidden-to-hidden connections, hence each cell/head can be computed independently of the others) allows to reformulate mLSTM in a parallel form, in order to speed up training when the full sequence is available in advance (see Appendix A.3 of the [15] for further details). When composed into an architecture of stacked blocks, connected via residual connections of two types (with post-up projections when considering sLSTM—like Transformers—or with pre-up projections when considering mLSTM), the model is referred to as an eXtended LSTM (xLSTM). Overall, xLSTM have a $\mathcal{O}(L)$ computational and $\mathcal{O}(1)$ memory complexities. Additionally, mLSTMs, despite being computationally expensive due to the presence of matrix-based state, implements a parallel-computation form, while sLSTM is not parallelizable due to memory mixing.
Constraining the Recurrence Weights.   Exploding and vanishing gradients hampers the RNNs’ ability to learn long-term dependencies. A recent strategy to circumvent this issue and allow the stable propagation of signals over long time scales, is to constrain the hidden-to-hidden weight matrix to be orthogonal or unitary (i.e., an element of the orthogonal group—referred to as Unitary and Orthogonal RNNs—see [5] and references therein), which ensures that the eigenvalues have unit norm and the dynamics are stable. However, despite advantages in terms of long-term memory preservation, this also reduces the expressivity of the model, as orthogonal transformations are limited in variety. We point the readers’ attention towards [253] (survey) and the extensive descriptions of several recent works, e.g., [67, 154]. Some recent works have proposed alternative ways to overcome this trade-off, such as using non-normal matrices with unit norm eigenvalues without orthogonality constraints on eigenbases (nnRNN [140]), or formulating the recurrent units by differential equations and updating the hidden states exploiting the difference between state values [131] These methods aim to improve the performance and flexibility of RNNs while preserving their long-term memory, without explicit constraints on the weight matrices.
@@ 5.1 ODE-inspired Recurrent Neural Networks
A recent trend involves recurrent architectures whose processing scheme is formalized by Ordinary Differential Equations (ODEs) in the context of dynamical systems. Two main branches of scientific works developed, the former based on continuous-time RNNs and the latter on discretized ODEs.
Continuous-time RNNs.   Continuous-time recurrent networks have been the subject of investigation since the dawn of neural networks and, later on, they were deeply investigated at the intersection of machine learning and other scientific fields, such as signal processing [178]. Amongst others, more recently, the interest on continuous-time RNNs has been renewed by studies on Neural ODEs [38] and ODE RNNs [245], where a continuous ODE acts as the learning model and gradients are computed from a sensitivity equation, which allows one to trade accuracy with computing time. The state of a Neural ODEs is defined by the solutions of the equation $\dot{x}=f(x,u,t,\theta)$, where $t$ represents continuous time, $u\vcentcolon=u(t)\in\mathbb{R}^{d_{\text{in}}}$ the time-dependent input signal, $x\vcentcolon=x(t)\in\mathbb{R}^{d_{\text{s}}}$ the RNN hidden state, $\dot{x}$ its first order time derivative and $f$ a neural network parametrized by $\theta$. Readers can find further details in [141] and references therein. Liquid Time-constant Networks [99], rather than defining the derivatives of the hidden-state directly by a neural network $f$ as in Neural ODE, determine a more stable continuous-time RNN in the form,
@@
where $\bar{A},\bar{W}$ are tunable matrices with an ad-hoc fixed structure. By leveraging tools from nonlinear systems theory, the authors carried on a stability analysis on the proposed recurrent unit behaviour in the long-term, resulting in good performances and expressivity. Recently, an in-depth analysis on approximation properties and optimization dynamics of continuous-time RNNs has been carried on in [161, 160], with an interesting take on the interaction of memory and recurrent structures in the linear dynamical setting.
@@
When discretizing such system with a fixed timestep $10<\Delta t<$ 1, the RNN hidden state at time $t_{n}=n\Delta t\in[0,1]$ evolves accordingly to the following laws,
@@
$\hskip 14.22636pt-\Delta t\gamma x_{n-1}-\Delta t\epsilon z_{\bar{n}},$
@@
with $\bar{n}$ either $\bar{n}=n$ or $\bar{n}=n-1$, depending on the fact that the damping term $\epsilon z$ is treated implicitly (the former case) or explicitly (the latter). In the coupled networks defined by Eq. (62), each neuron updates its hidden state based on input signals and information from other neurons. The diagonal entries of $W$ and the hyperparameter $\gamma$ control oscillation frequency, while the diagonal entries of $\hat{W}$ and the hyperparameter $\epsilon$ determine damping for each neuron, whereas non-diagonal entries modulate interactions between neurons. Input signals drive the generation of (superpositions of) oscillatory wave-forms, controlled by the other tunable parameters. This leads to rich global dynamics and the emergence of non-trivial non-oscillatory hidden states from oscillatory inputs, emphasizing the network’s high expressivity in approximating outputs from complex sequential inputs. The authors derive bounded gradients and limited hidden state magnitude for the coRNN model, under some mild assumptions. Thus, coRNN has stable dynamics which foster better performances than existing RNNs, especially on tasks with very long time-dependencies. In this family of models, which is recently referred to as Neural Oscillators [149], lies UnICORNN [249], a multi-layer sequence model that stacks networks of independent (uncoupled) undamped oscillators as hidden layers within an RNN. In contrast to coRNN, neurons in UnICORNN are independent (uncoupled) and as there is no damping, the ODE system yielding UnICORNN has an Hamiltonian structure. This characteristic allows the model to avoid any assumptions on the weights, whereas the mitigation of exploding/vanishing gradients in coRNN was depending on specific restrictions imposed on the weights. Moreover, Neural Oscillators have been proven to be capable to approximate any continuous and casual operator mapping between time-varying functions, to desired accuracy [149]. Their performances on long-range sequences are remarkable [249]. Locally coupled oscillatory recurrent neural network has been used to model the neuroscience concept of traveling waves, referred to as Neural Wave Machines (NWMs) [138]. Such waves serve as a bias towards learning structured representations, which exhibit complex spatio-temporal dynamics when modeling real data. When tasked to reconstruct the input signal, NWMs use traveling waves to encode transformations in the RNN hidden state. Waves-like dynamics can be modeled also with simpler RNN architectures though connectivity constraints and initialization [137], and can act as memory storage system on complex sequence modeling. NoisyRNN [164] consider discretizations of the stochastic differential equations (SDEs) obtained from ODE formulations of RNNs through the addition of a diffusion (noise) term, as an implicit regularization. By dropping the noisy elements at inference time, NoisyRNN can be considered as a stochastic learning strategy (i.e., similarly to Dropout) with several advantages such as more stable dynamics. This introduces a form of implicit regularization leading towards the development of classifiers with a large classification margin, that keep generalization error small. However, despite the stabilization properties, noise injection could negatively impact capacity for long-term memory. Rusch et al. [250] pointed out that real-world data could contain information arranged according to multiple scales, i.e., time, lengths etc., depending on the considered data and task. They propose Long Expressive Memory (LEM), based on a time-discretization of a set of multiscale ODEs. These scales can be learned adaptively (with respect to states) and dynamically (in time). LEM has bounded gradients that mitigate exploding/vanishing issues, and favour the model ability in the context of long sequence processing. Irie et al. [116] introduced learning rules and Neural ODEs to build continuous-time sequence processing nets that learn to manipulate short-term memory in rapidly changing synaptic connections of other nets. This yields continuous-time counterparts of Fast Weight Programmers and Linear Transformers [118, 119]. Learning rules can be seen as the outcome of discretization procedures applied to ODEs. Kag et al. [131] proposed a modified differential equation for the state update, obtained by leveraging implicit discretization methods instead of the (explicit) Euler method, to foster the system stability. The main intuition is that the hidden states are updated based on the difference between predicted and previous states. Then, the implicit equation is solved via fixed-point recursion, with stable fixed-points and fast convergence. In a subsequent work, Kag et al. [129] also explored a time-adaptive discretization of the ODE where time-steps are modified in function of the current observation and the hidden state.

### v2

@@ 5 Enhancing Recurrent Neural Networks
This section gathers recent approaches that are not directly related to the previous macro-categories of Transformer architectures and State-Space Models, but still focus on improving recurrent models. It turns out that several of the architectural trends described in the previous sections (i.e., element-wise linear recurrence, novel gating mechanisms, etc.), and other new ones, are also significantly explored in the scientific literature that aims to address two of the main drawbacks of RNN , namely slow sequential training and limited capability in modeling long-term dependencies. In Figure 16, we report an overview of the main topics/approaches covered by this section.
Figure 16: Conceptual overview of the organization of Section 5 , which categorizes novel recurrent models based on their architectural innovations and modeling strategies. Blue boxes indicate subsections, yellow boxes denote thematic categories, and representative models are listed in gray.
Simplifying RNNs to Gain Speed.   RNNs are based on sequential processing of the input data, which does not directly allow for the development of efficient implementations that process the input tokens in parallel or update the components of the hidden state in parallel. It turns out that this limit is mostly due to ($i$) the non-linearity applied to recurrent layers, and ($ii$) the fact that updates in the hidden state are performed by full matrix multiplication, due to a dependency on all components of the hidden state from the previous time step. In detail, the standard update scheme of RNNs (Eq. 2) assumes that all the neurons in one layer contribute to the state computation of every other neuron (i.e., through the $Ax_{t-1}$ term). Each element of the state vector $x_{t}$ depends on all the entries of $x_{t-1}$. Early works, such as Independently RNN (IndRNN) [31], propose layers composed of “independent” neurons, achieved by modifying Eq. (2) as,
@@
where Eqs. (57)-(58) represent the proposed “lightweight” recurrence, where the reset gate $r_{t}$ adaptively combines the input and the cell state $c_{t}$, with learnable vectors $a_{f},a_{r},b_{r},b_{f}$ and learnable matrices $B_{f},B_{r}$. The skip connection in Eq. (60) favours gradient propagation. Independence between distinct hidden states enables efficient element-wise product instead of full matrix multiplication (i.e., in classical forget gates dense matrices products inject a dependencies on all neurons previous states), as the authors show when the (nonlinear) recurrence is fused within a single CUDA kernel. This allows to reduce the complexity to $\mathcal{O}(Lb{d_{\text{in}}})$, where $b$ denotes here the batch dimension, while a standard LSTM takes $\mathcal{O}(Lb{d_{\text{in}}}^{2})$. The seminal work by Martin & Cundy [152] highlights that linear recurrences in the form of $x_{t}=\Lambda_{t}x_{t-1}+u_{t}$ are specific instances of the scan operation, a computation involving the repeated application of a binary operator over an array of data. This allows for a highly parallelizable unrolling of the recurrence using parallel scans [165], resulting in substantial improvements in training speeds. When $\Lambda_{t}$ is diagonal, the cost of a ELR with parallel scan and $p$ processors is $\mathcal{O}\left(6{d_{\text{s}}}(L/p+\text{log}p)\right)$, while the cost of serial scan is $\mathcal{O}(2{d_{\text{s}}}L)$. This reduction becomes important when the sequence length $L$ is large since, given sufficient processors $p$, the parallel time scales logarithmically with the sequence length. Please refer to Appendix H of [42] for a detailed overview of the parallel scan operation. Moreover, several existing models (QRNN[162], SRU [164]) fall under this class of approaches, and the authors of [152] provide an efficient CUDA kernel that speed up their training. This work laid the foundations for the efficient parallel implementation of SSMs such as S5 [42] and others [43] (Section 4). Additionally, while typical forget gate values depend on both the previous hidden state and the current input, the authors suggest that forget gate values should depend solely on the current inputs to enable parallel training. Lately, Lim et al. [166] built on top of [152] and have shown that it is also possible to parallelize the evaluation and training of non-linear sequential models like classic RNN and NeuralODE [167], by introducing a general framework to solve non-linear differential equations, which are restated as fixed-point iteration problems with quadratic convergence, equivalent to Newton’s method. Each fixed-point iteration involves parallelizable operations and an inverse linear operator that can be evaluated in parallel even for the aforementioned sequential models, resulting in improvements of up to $3$ orders of magnitude in terms of speed in the case of long sequences.
@@
Enhancing Gating Mechanisms.   Learning long-term dependencies requires the ability to modulate the effect of the incoming information. Several recent studies have incorporated gating mechanisms inspired by LSTMs [28] (or related intuitions) into SSMs, which are characterized by gates acting on linear recurrence layers [168, 146, 42], yielding impressive performance gains.
Gu et al. [168] investigated the saturation regime of gates, remarking the fact that capturing long-term dependencies in gated RNNs requires forget gate values close to one. Unfortunately, learning with gates in their saturated regimes (i.e., values close to 0 or $1$) is difficult, since they suffer from vanishing gradients. Moreover, if all forget gate values are close to one, the model’s ability to forget irrelevant information is hindered. To overcome such issues, the authors of [168] propose to tweak the forget gate $f_{t}$ with an independent refine gate $r_{t}$ that is exploited to produce an input-dependent bounded additive term $\phi(f_{t},u_{t})$ that modulates $f_{t}$, allowing much higher/lower activations,
@@
where $W_{r}^{*}$ denote projection matrices and $\hat{c}_{t}$ is the cell input activation vector from LSTMs. The form of the additive update $\phi(f_{t},u)$ emerges from the consideration that gating mechanisms should be bounded in $[0,1]$, symmetric around 0, and differentiable. Additionally, the authors propose to initialize the forget gate $f_{t}$ with values sampled from the uniform distribution $\mathcal{U}(0,1)$, instead of constant values, even allowing negative biases. This choice fosters the gate’s ability to grasp different timescales, as noticed in the chrono initialization approach by Tallec & Olivier [169].
In previous Sections, we showed how several linear RNNs with static decay rates perform eigendecompositions on the recurrent weight matrix to achieve element-wise linear recurrence [114, 43]. Notice that, if only real-valued eigenvalues are allowed, this choice restricts the range of the recurrent weight matrix to be symmetric, limiting the expressiveness of the model. To overcome this limitation, linear RNNs often employ complex-valued eigenvalues, [40, 43]. Following such intuitions, Hierarchically Gated Recurrent Units (HGRU) [131] exploit linear recurrence in the complex domain, and address the saturating issue pointed out by Gu et al. [168] by adding an additive learnable value $\Gamma$ to the original forget gate, with the purpose of pushing gate activations away from the saturated regimes. The $\Gamma$ variable, which acts as a lower bound on forget gate values, is forced to increase monotonically with the model depth, inspired from Ordered Neuron LSTM [170]: small value in lower layers, in order to ease the forgetting of past-information (short-term dependencies); forget value close to one in top-most layers, facilitating the modeling of long-term dependencies. In detail, HGRU leverages a gated linear recurrent as follows,
@@
where $\Gamma\in\mathbb{R}^{H\times{d_{\text{s}}}}$ independently parametrizes the lower bounds for all hidden states, where cumsum and softmax operate over the first dimension of their tensor input. Basically, the composition of the the softmax and cumsum functions forms an activation function which yields a monotonically increasing vector in $[0,1]$. The squared bracket-based notation is defined as $\text{cumsum}(x)]_{l}=(\sum_{i=1}^{l}x_{i})-x_{1}$. Notice that $\text{cumsum}_{0}$ is applied to the layer dimension across different layers, to enable upper layers to model long-range dependencies. Then, the model exploits an output projection with a learned gate, similarly to what happens in SSMs. Another perspective on the role of gating mechanisms can be appreciated by inspecting the already described connection between the update mechanisms in linear RNNs and linear Transformers, discussed in previous Sections. Recently, Zucchet et al. [171] proposed a unifying view on such architectures, driven by the role of gating functions. In particular, under a specific parametrization (that leverages GLU [134] and requires a number of parameters squared with respect to attention parameters), they showed that RNNs equipped with linear recurrent layers interconnected by feed-forward paths with multiplicative gating can, through learning, encode attention-based algorithms disguised in their weights.
@@
where $q_{t},k_{t},v_{t}\in\mathbb{R}^{d_{\text{k}}}$ are linear projections of the input (such as in self-attention), while $w_{i},w_{f}\in\mathbb{R}^{{d_{\text{in}}}}$ are learnable weight vectors. In the cell-state update rule, the forget gate acts as a decay rate, the input gate as a learning rate, while the output gate scales the vector which is retrieved by the outer product. The normalizer state $n_{t}$, which keeps a record of the “strength” of the gates, is the weighted sum of the key vectors, where, in turn, each key is weighted by the input gate and the current forget gate. In mLSTM, considering multiple cells is equivalent to multiple heads, since in this case there is no memory mixing, due to the presence of matrix states. Interestingly, the absence of memory mixing (no hidden-to-hidden connections, hence each cell/head can be computed independently of the others) allows us to reformulate mLSTM in a parallel form, in order to speed up training when the full sequence is available in advance (see Appendix A.3 of [172] for further details). When composed into an architecture of stacked blocks, connected via residual connections of two types (with post-up projections when considering sLSTM—like Transformers—or with pre-up projections when considering mLSTM), the model is referred to as an eXtended LSTM (xLSTM). Overall, xLSTM have a $\mathcal{O}(L)$ computational and $\mathcal{O}(1)$ memory complexities. Additionally, mLSTMs, despite being computationally expensive due to the presence of matrix-based state, implements a parallel-computation form, while sLSTM is not parallelizable due to memory mixing.
Constraining the Recurrence Weights.   Exploding and vanishing gradients hamper the RNNs’ ability to learn long-term dependencies. A recent strategy to circumvent this issue and allow the stable propagation of signals over long time scales, is to constrain the hidden-to-hidden weight matrix to be orthogonal or unitary (i.e., an element of the orthogonal group—referred to as Unitary and Orthogonal RNNs—see [174] and references therein), which ensures that the eigenvalues have unit norm and the dynamics are stable. However, despite advantages in terms of long-term memory preservation, this also reduces the expressivity of the model, as orthogonal transformations are limited in variety. We draw readers’ attention to [18] (survey) and the extensive descriptions of several recent works, e.g., [175, 176]. Some recent works have proposed alternative ways to overcome this trade-off, such as using non-normal matrices with unit norm eigenvalues without orthogonality constraints on eigenbases (nnRNN [177]), or formulating the recurrent units by differential equations and updating the hidden states exploiting the difference between state values [178] These methods aim to improve the performance and flexibility of RNNs while preserving their long-term memory, without explicit constraints on the weight matrices.
@@ 5.1 ODE-inspired Recurrent Neural Networks
A recent trend involves recurrent architectures whose processing scheme is formalized by Ordinary Differential Equations (ODEs) in the context of dynamical systems. Two main branches of scientific works have developed, the former based on continuous-time RNNs and the latter on discretized ODEs.
Continuous-time RNNs.   Continuous-time recurrent networks have been the subject of investigation since the dawn of neural networks and, later on, they were deeply investigated at the intersection of machine learning and other scientific fields, such as signal processing [179]. Amongst others, more recently, the interest in continuous-time RNNs has been renewed by studies on Neural ODEs [167] and ODE RNNs [180], where a continuous ODE acts as the learning model and gradients are computed from a sensitivity equation, which allows one to trade accuracy with computing time. The state of a Neural ODEs is defined by the solutions of the equation $\dot{x}=f(x,u,t,\theta)$, where $t$ represents continuous time, $u\vcentcolon=u(t)\in\mathbb{R}^{d_{\text{in}}}$ the time-dependent input signal, $x\vcentcolon=x(t)\in\mathbb{R}^{d_{\text{s}}}$ the RNN hidden state, $\dot{x}$ its first order time derivative and $f$ a neural network parametrized by $\theta$. Readers can find further details in [181] and references therein. Liquid Time-constant Networks [153], rather than defining the derivatives of the hidden-state directly by a neural network $f$ as in Neural ODE, determine a more stable continuous-time RNN in the form,
@@
where $\bar{A},\bar{W}$ are tunable matrices with an ad-hoc fixed structure. By leveraging tools from nonlinear systems theory, the authors carried out a stability analysis on the proposed recurrent unit behaviour in the long-term, resulting in good performance and expressivity. Recently, an in-depth analysis of approximation properties and optimization dynamics of continuous-time RNNs has been carried out in [183, 64], with an interesting take on the interaction of memory and recurrent structures in the linear dynamical setting.
@@
When discretizing such system with a fixed timestep $0<\Delta t<1$, the RNN hidden state at time $t_{n}=n\Delta t\in[0,1]$ evolves accordingly to the following laws,
@@
$\qquad-\Delta t\gamma x_{n-1}-\Delta t\epsilon z_{\bar{n}},$
@@
with $\bar{n}$ either $\bar{n}=n$ or $\bar{n}=n-1$, depending on the fact that the damping term $\epsilon z$ is treated implicitly (the former case) or explicitly (the latter). In the coupled networks defined by Eq. (62), each neuron updates its hidden state based on input signals and information from other neurons.The diagonal entries of $W$ and the hyperparameter $\gamma$ control oscillation frequency, while the diagonal entries of $\hat{W}$ and the hyperparameter $\epsilon$ determine damping for each neuron, whereas non-diagonal entries modulate interactions between neurons. Input signals drive the generation of (superpositions of) oscillatory wave-forms, controlled by the other tunable parameters. This leads to rich global dynamics and the emergence of non-trivial non-oscillatory hidden states from oscillatory inputs, emphasizing the network’s high expressivity in approximating outputs from complex sequential inputs. The authors derive bounded gradients and limited hidden state magnitude for the coRNN model, under some mild assumptions. Thus, coRNN has stable dynamics which foster better performance than existing RNNs, especially on tasks with very long time-dependencies. In this family of models, which is recently referred to as Neural Oscillators [185], lies UnICORNN [186], a multi-layer sequence model that stacks networks of independent (uncoupled) undamped oscillators as hidden layers within an RNN. In contrast to coRNN, neurons in UnICORNN are independent (uncoupled) and as there is no damping, the ODE system yielding UnICORNN has a Hamiltonian structure. This characteristic allows the model to avoid any assumptions on the weights, whereas the mitigation of exploding/vanishing gradients in coRNN was dependent on specific restrictions imposed on the weights. Moreover, Neural Oscillators have been proven to be capable of approximating any continuous and causal operator mapping between time-varying functions, to desired accuracy [185]. Their performances on long-range sequences are remarkable [186]. Locally coupled oscillatory recurrent neural networks have been used to model the neuroscience concept of traveling waves, referred to as Neural Wave Machines (NWMs) [187]. Such waves serve as a bias towards learning structured representations, which exhibit complex spatio-temporal dynamics when modeling real data. When tasked to reconstruct the input signal, NWMs use traveling waves to encode transformations in the RNN hidden state. Waves-like dynamics can be modeled also with simpler RNN architectures though connectivity constraints and initialization [188], and can act as memory storage system on complex sequence modeling. NoisyRNN [189] consider discretizations of the stochastic differential equations (SDEs) obtained from ODE formulations of RNNs through the addition of a diffusion (noise) term, as an implicit regularization. By dropping the noisy elements at inference time, NoisyRNN can be considered as a stochastic learning strategy (i.e., similarly to Dropout) with several advantages such as more stable dynamics. This introduces a form of implicit regularization leading towards the development of classifiers with a large classification margin, that keep generalization error small. However, despite the stabilization properties, noise injection could negatively impact capacity for long-term memory. Rusch et al. [190] pointed out that real-world data could contain information arranged according to multiple scales, i.e., time, lengths etc., depending on the considered data and task. They propose Long Expressive Memory (LEM), based on a time-discretization of a set of multiscale ODEs. These scales can be learned adaptively (with respect to states) and dynamically (in time). LEM has bounded gradients that mitigate exploding/vanishing issues, and favour the model ability in the context of long sequence processing. Irie et al. [191] introduced learning rules and Neural ODEs to build continuous-time sequence processing nets that learn to manipulate short-term memory in rapidly changing synaptic connections of other nets. This yields continuous-time counterparts of Fast Weight Programmers and Linear Transformers [94, 96]. Learning rules can be seen as the outcome of discretization procedures applied to ODEs. Kag et al. [178] proposed a modified differential equation for the state update, obtained by leveraging implicit discretization methods rather than the (explicit) Euler method, to foster system stability. The main intuition is that the hidden states are updated based on the difference between predicted and previous states. Then, the implicit equation is solved via fixed-point recursion, resulting in stable fixed points and fast convergence. In a subsequent work, Kag et al. [192] also explored a time-adaptive discretization of the ODE where time steps are modified based on the current observation and the hidden state.
This section catalogued architectural innovations that aim to simplify, stabilise or extend classic RNNs: independent/diagonal updates for parallelism, unitary or orthogonal weights for gradient preservation, enhanced gating (e.g., xLSTM), and continuous-time formulations such as Neural ODEs. In the next Section 6, we summarize recently introduced efficient and forward-in-time alternatives to BPTT used to train recurrent models.